OpenStudy (anonymous):

A cleans 40 plates in 20 minute. B cleans 35 plates in 30 minute. How long would it take to clean 60 plates if they were to work together?

OpenStudy (anonymous):

i hate word problems, don't you?

OpenStudy (anonymous):

doesn't mean we can't do it though

OpenStudy (anonymous):

Its going to pop up on the SAT, gotta learn how to do these problems!

OpenStudy (anonymous):

lol there is motivation for you

OpenStudy (anonymous):

A does 40 plates in 20 minutes, so his rate per minute is \(\frac{40}{20}=2\) plates per minute

OpenStudy (anonymous):

\[b = \frac{ 35 }{ 30 }\] or 1.16 plates per minute

OpenStudy (anonymous):

B is slower, 35 plates in 30 minutes that rate is \(\frac{35}{30}=\frac{7}{6}\)

OpenStudy (anonymous):

lets not use decimals

OpenStudy (anonymous):

combined rate is what you get when you add them

OpenStudy (anonymous):

I'll just leave this one to you ;)

OpenStudy (anonymous):

\[2+\frac{7}{6}=\frac{19}{6}\] if i still know how to add

OpenStudy (anonymous):

you want the time, set \[\frac{19}{6}=60\] and solve for \(t\) via \[t=60\times \frac{6}{19}\]

OpenStudy (anonymous):

wow typo there

OpenStudy (anonymous):

set \[\frac{19}{6}T=60\] get \[T=60\times \frac{6}{19}\]

OpenStudy (anonymous):

did you cross multiplied?

OpenStudy (anonymous):

and 19x1 and 6x60?